OpenStudy (anonymous):

Will Fan And Give Medal What is the solution to the system? Line 1 x y –2 –4 –1 –1 Line 2 x y 2 3 3 1

OpenStudy (aripotta):

well you wanna get the equations of the lines first. so to do that, we're gonna get the slope of line 1 first. so use the two points (-2, -4) and (-1, -1) and put them in the slope formula \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

it says this

OpenStudy (aripotta):

yes i know the question. but you need the equations of the lines first

OpenStudy (aripotta):

so what's the slope of line 1?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

Well, looks like u have a Mathlete helping you .-.

OpenStudy (aripotta):

from the table, we see two points on line 1, (-2, -4) and (-1, -1). use those points and plug them into the slope formula, which is \[\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

1,2 @AriPotta

OpenStudy (aripotta):

\[\frac{ -4-(-1) }{ -2-(-1) }=~\frac{ -4+1 }{ -2 + 1 }=~\frac{ -3 }{ -1 }=~3\]so the slope is 3

OpenStudy (anonymous):

oh ok but whats the answer then

OpenStudy (aripotta):

i'm not giving you the answer

OpenStudy (aripotta):

now that we know the slope is 3, we're going to use that and one of the points into point-slope form, which is y - y1 = m(x - x1)

OpenStudy (anonymous):

so its 3,1

OpenStudy (anonymous):

@AriPotta

OpenStudy (anonymous):

@Angel_Bear19

OpenStudy (anonymous):

@Angel_Bear19 can u help

OpenStudy (anonymous):

idk sorry, @iGreen may be on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@iGreen